# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v1

@@ 5 Conclusion and future work
In this paper, we introduce a structural component within LLM-based agents named the core-agent. This component is engineered to address the architectural ambiguities that software developers encounter and enhance modularity. We propose LLM-Agent-UMF, a comprehensive framework for modeling the structure of the agent, explicating each of core-agent’s five modules: planning, memory, profile, action, and security. Subsequently, we classified core-agents into passive and active and highlighted their structural and functional differences. Based on this classification, we designed uniform and hybrid multi-core agent architectures. Most prominently, the one-active-many-passive architecture exploits the full potential of both active and passive core-agents, striking a balance between easiness of development and the power of hybrid architectures. By applying our framework to state-of-the-art agents, we identified within their structures core-agents and their constituting internal modules which assisted us in the classification process. This allowed us to recognize the individual characteristics of each agent and discover potential prospects of merging different functionalities into a single multi-core agent.
@@
In conclusion, the ultimate purpose of this framework, which is predicated on the core-agent, is to reformulate the conception of LLM-based agents. By basing their development on this unit rather than addressing it in a monolithic manner, researchers and practitioners can use a unified terminology to refer to different modules within a common architecture. This shared foundation facilitates uniformity in research and enhancements. Additionally, developers can commence implementing consistent solutions that are easily maintainable and highly adaptable for future improvements.
Appendix A Screenshot of ChatGPT 4o mini refusing to explain how to hot-wire a car
Figure 14: ChatGPT 4o mini refusing to explain how to hot-wire a car

### v2

@@ 5 Conclusion and future work
In this paper, we introduce a structural component within LLM-based agents named the "core-agent". This component is engineered to address the architectural ambiguities that software developers encounter and foster better understanding of the interacting entities within LLM-powered agents. We propose the LLM-Agent-UMF, a comprehensive framework for modeling the structure of the agent, explicating each of core-agent’s five modules: planning, memory, profile, action, and security. Subsequently, we classified core-agents into passive and active categories and highlighted their structural and functional differences. Based on this classification, we designed uniform and hybrid multi-core agent architectures. Most prominently, the one-active-many-passive architecture exploits the full potential of both active and passive core-agents, striking a balance between easiness of development and the power of hybrid architectures. By applying our framework to state-of-the-art agents, we identified within their structures core-agents and their constituting internal modules which assisted us in the classification process. This allowed us to recognize the individual characteristics of each agent, discern their limitations and discover potential prospects of merging different functionalities into a single multi-core agent.
@@
In conclusion, the ultimate purpose of this framework, which is predicated on the core-agent, is to reformulate the conception of LLM-based agents. By basing their development on this unit rather than addressing it in a monolithic manner, researchers and practitioners can use a unified terminology to refer to different modules within a common architecture. This shared foundation not only facilitates uniformity in research and development but also enables developers to implement consistent solutions that are easily maintainable and highly adaptable for future improvements.
CRediT authorship contribution statement
Amine Ben Hassouna: Writing – original draft, Writing – review & editing, Methodology, Project administration, Supervision, Validation, Visualization, Investigation, Data Curation, Formal analysis, Conceptualization. Hana Chaari: Writing – original draft, Methodology, Investigation, Data Curation, Formal analysis. Ines Belhaj: Writing – original draft, Methodology, Investigation, Data Curation, Formal analysis.